OpenStudy (anonymous):

Outcome Payout HHH $10 HHT $3 HTH $0 HTT -$3 TTT $10 THH -$3 THT $0 TTH $0 Kara is playing a game where she flips a coin 3 times. She wins and loses different amounts of money based on the outcomes. What is the EXPECTED VALUE for the game?

OpenStudy (anonymous):

multiply the amount you win times the probability you win that amount and add it all up

OpenStudy (anonymous):

these are the answer choices A) $0 B) $2.13 C) $3.29 D) $4.14 but i keep getting like 1.25

OpenStudy (anonymous):

you can ignore the payoffs of zero, because they contribute nothing, so it will be \[\frac{10+3-3-3}{8}\]

OpenStudy (anonymous):

no that was wrong \[\frac{10+10+3-3-3}{8}\]

OpenStudy (anonymous):

look like \[\frac{17}{8}=2.125\] so i guess you are supposed to go with $2.13

OpenStudy (anonymous):

2.125? so round to 2.13?

OpenStudy (anonymous):

yeah i think so, round since it is money but really it is 2.125 right

OpenStudy (anonymous):

so is it like just the average?

OpenStudy (anonymous):

exactly like an average, yes

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

yw